# Extreme Nuclear Transients Resulting from the Tidal Disruption of Intermediate Mass Stars

### v1

@@
Extreme Nuclear Transients Resulting from the Tidal Disruption of Intermediate Mass Stars
Modern transient surveys now routinely discover flares resulting from tidal disruption events (TDEs) which occur when stars, typically M⊙, are ripped apart after passing too close to a supermassive black hole. We present three examples of a new class of extreme nuclear transients (ENTs) that we interpret as the tidal disruption of intermediate mass ( M⊙) stars. Each is coincident with their host-galaxy nucleus and exhibits a smooth (% excess variability), luminous ( erg s-1), and long-lived ( days) flare. ENTs are extremely rare ( Gpc-1 yr-1) compared to any other known class of transients. They are at least twice as energetic ( erg) as any other known transient and these extreme energetics rule out stellar origins.
@@
Accretion onto supermassive black holes (SMBHs) powers many of the most luminous events in the universe. At a redshift of , roughly 10% of SMBHs are actively accreting mass (e.g., (?)) and are observed as active galactic nuclei (AGNs). AGN light curves commonly show stochastic variability at a broad range of timescales from minutes to years (e.g., (?, ?)), with some AGNs showing long-term photometric trends often accompanied by dramatic changes in their spectra (e.g., (?, ?, ?)). AGNs can also, albeit rarely, exhibit large, coherent flares (?), although the physical mechanisms for powering them are unclear.
@@
With the recent growth of optical transient surveys, several classes of flares coincident with the nuclei of their host galaxies have been detected. These include tidal disruption events (TDEs; (?, ?, ?)), rapid turn-on AGNs (?, ?), and ambiguous nuclear transients (ANTs; (?, ?, ?)). Accretion-powered transients share several key observational properties including bright ultraviolet (UV) emission (?, ?, ?), strong emission lines (?, ?, ?), and often X-ray emission (?, ?, ?, ?). The smooth flares of nuclear transients on several-month timescales (?, ?) are distinct from the stochastic variability typical of AGNs.
@@
A TDE results from the disruption of a star as it passes too close to a SMBH (e.g., (?, ?)). Most observed TDEs appear consistent with the disruption of a main-sequence star with a mass of M⊙ (?), although there is significant scatter in these estimates (?). Nevertheless, characteristics like enhanced N/C ratios (?, ?, ?) suggest a population of TDEs resulting from more massive stars. The host galaxies of TDEs typically do not host a strong AGN, although this is largely a result of selection effects. Recently, an increasing number of TDE candidates have been discovered for which their host galaxies exhibit signs of weak AGN activity (e.g., (?, ?, ?)).
@@
Owing to the establishment of long-baseline all-sky surveys (?, ?, ?), we are now sensitive to rare and unexpected classes of transients. One such survey, Gaia Alerts, uses the Gaia spacecraft to monitor the transient sky at approximately monthly cadence with high precision. From the Gaia Alerts (?) transient stream we selected a sample of flares with three primary characteristics: (1) large amplitudes of 1 mag, (2) smooth light curves with 10% excess variability about the flare evolution, and (3) a long timescale of year. Gaia is ideal for such a search as it has observed the full sky since late 2014 and, as a space-based mission, it typically has shorter seasonal breaks than ground-based surveys. Our search yielded two transients, Gaia16aaw (AT2016dbs) and Gaia18cdj (AT2018fbb). We combine these events with the recently published object ZTF20abrbeie (AT2021lwx; (?, ?)) as a sample of events we will refer to as extreme nuclear transients (ENTs).
@@
The observed properties of the ENTs are reminiscent of extreme versions of ANTs, which are transients occurring in an AGN host galaxy. The light curves of the ENTs, shown in Figure 1, each exhibit a long ( 100 day) rise to a high peak luminosity. The ENTs decline slowly after peak, taking more than 150 days to fade to half of their peak luminosity. The ENTs detected prior to the flare show tentative signs of pre-flare variability, suggesting weak AGN activity within their host galaxies. After the UV/optical emission peaks, the ENTs show an IR excess, indicative of transient heating of circumnuclear dust and re-emission at longer wavelengths. Much like the ENT hosts, the host galaxies of AGNs typically have significant nuclear dust.
@@
The ENTs Gaia16aaw and Gaia18cdj are located within 0.8 kpc of their host-galaxy centers, confirming that they are nuclear transients. Their nuclear origin, long timescales, and high peak luminosities are immediately suggestive of a transient resulting from accretion onto a SMBH. We cannot measure the host offset of AT2021lwx as it has no detected host galaxy prior to the flare, at levels of absolute mag in the rest-frame blue bands. Nevertheless, the similar timescales and peak luminosities of AT2021lwx indicate that it is also powered by accretion, consistent with previous studies.
@@
The ENTs are located at a relatively high redshift of , as measured from optical and near-IR follow-up spectra. Using the stellar population synthesis and AGN model of the Code Investigating GALaxy Emission (CIGALE; (?)), we find that the host galaxies of Gaia16aaw and Gaia18cdj each have a stellar mass of M⊙ and star-formation rates (SFRs) of M⊙ yr-1. While undetected, the luminosity limits for the host galaxy of AT2021lwx combined with a conservative mass-to-light ratio of 3 yields a mass upper-limit of M⊙. The upper limit on the [O II] emission means that the SFR for AT2021lwx is M⊙ yr-1.
@@
From typical galaxy-SMBH scaling relations (?), the stellar masses imply SMBH masses of M⊙ for Gaia16aaw and Gaia18cdj and an upper limit on the mass of M⊙ for AT2021lwx, which are more massive SMBHs than in the majority of known nuclear transient hosts. The detected ENT hosts are more massive and display higher SFRs than the host galaxies of local nuclear transients such as TDEs and ANTs. Furthermore, the detected ENT host-galaxy masses are within the top few percent of stellar masses at , when the universe was half its current age. Thus, the inferred SMBH masses are similarly extreme at this redshift. In contrast, the prodigious SFRs of the hosts of Gaia16aaw and Gaia18cdj are only moderately high at when the star formation density in the universe was a factor of 6 higher than today (?).
@@
The rest-frame spectra of the ENTs, shown in Figure 2, exhibit blue spectra with broad lines from the Balmer series of hydrogen and singly-ionized magnesium (Mg II). These are similar to the comparison spectra of the luminous nuclear transients ASASSN-15lh (?, ?), AT2019brs (?), ASASSN-17jz (?), PS1-10adi (?), ASASSN-18jd (?), PS16dtm (?), and AT2019dsg (?) also shown in Fig. 2. The persistent blue continua and broad lines are inconsistent with known classes of supernovae, but fully consistent with SMBH accretion. The Mg II and H emission of the ENTs is broad ( km s-1) and luminous ( erg s-1), very similar to AGNs. While Mg II emission has not been seen for TDEs, the H emission is consistent with TDEs if the lines are equally as overluminous as the broadband emission.
As the ENTs are likely powered by accretion, it is important to consider the presence of previous AGN activity. Through a combination of the CIGALE fits, WISE IR colors, narrow emission lines, and X-ray emission we find evidence of a strong AGN in Gaia16aaw. Gaia18cdj likely hosts a weak AGN, but does not host a strong quasar. The limits derived from the pre-flare properties of AT2021lwx are consistent with a weak AGN, and rule out a strong quasar. The fact that some ENT host galaxies do not host a strong AGN suggests that prior strong AGN activity is not a requirement to power an ENT.
@@
The ENT spectral energy distributions (SEDs) are well-fit by a blackbody model, which is consistent with the super-Eddington accretion expected from TDEs. In contrast, AGNs with broad lines typically exhibit power-law-like SEDs as a result of viewing the accretion disk directly. The resulting bolometric light curves are shown in Figure 3, along with the comparison objects from Fig. 2. The ENTs are extremely luminous, with peak luminosities from erg s-1. This is times more luminous than typical core-collapse supernovae (SNe), times the average Type Ia SN peak luminosity, and times more luminous than typical SLSNe-I. Only ASASSN-15lh, suggested to either be a superluminous supernovae (SLSNe; (?)) or TDE (?), rivals these high peak luminosities.
@@
The blackbody properties of the ENTs are also consistent with some form of accretion onto a SMBH rather than an exotic class of SN. The temperatures of the ENTs are hot at K and show little or very slow evolution during the flare. This is inconsistent with SLSNe, which quickly cool as the ejecta expands. However, this behavior is similar to TDEs and ANTs, with a remarkable agreement between the ANT and ENT blackbody temperatures. The large effective radii of the ENTs is consistent with SLSNe and some ANTs, although the decreasing blackbody radii in time is more typical of TDEs and ANTs.
@@
The light curve decay timescale of the ENTs is also far longer than most transients, which is consistent with TDEs occurring on massive SMBHs. The rest-frame durations for the flares to fade by half are days for Gaia16aaw, days for Gaia18cdj, and days for AT2021lwx, respectively. Figure 4 shows the position of these ENTs in the parameter space of peak absolute magnitude and characteristic timescale. The ENTs stand out in this parameter space for being very luminous and long-lived. ASASSN-15lh, which has a similar peak luminosity to the ENTs, decays more quickly, with timescale of days. The position of the ENTs in the upper right corner of this space indicates a high total radiated energy.
@@
The radiated energies of the ENTs are erg for Gaia16aaw, erg for Gaia18cdj, and erg for AT2021lwx. These extreme radiated energies correspond to high accreted masses of M⊙ for a typical 10% accretion efficiency, far greater than typical TDEs and ANTs. The ENT flares are at least twice as energetic as the next most energetic known flares, PS1-10adi (?) and ASASSN-15lh (?, ?), and up to an order of magnitude higher in the cases of Gaia18cdj and AT2021lwx. While the estimates for accreted mass are extreme compared to local transients, they are fully consistent with the tidal disruption of an intermediate-mass (3-10 M⊙) star.
@@
The dust properties of ENTs are very similar to other SMBH accretion-powered transients. To study the environment directly surrounding the SMBHs, we used NEOWISE data to probe the emission from hot dust as it reprocesses the intense UV/optical emission from the flare. By fitting the WISE IR SEDs as blackbodies, we find peak luminosities of erg s-1, temperatures of K and radii of pc, all consistent with hot dust in nuclear environments (?, ?). From the ratio of the peak IR luminosity to the peak UV/optical luminosity we estimate dust covering fractions of , which we confirm with models of the optical and IR light curves (see the supplementary text). These covering fractions are consistent with AGN dust covering fractions at similar SMBH masses as well as dust-obscured TDE candidates (?). This indicates the presence of significant dense gas and dust near the SMBH, which likely supports the existence of AGN activity, whether weak or in the past, in each of the ENT hosts.
@@
Accretion-powered transients often show X-ray emission, as do two of the three ENTs in our sample. Gaia16aaw and AT2021lwx both show X-rays at levels of erg s-1 in the 0.3–10 keV band throughout the flare, similar to luminous AGNs (?). Gaia18cdj is undetected in the X-rays at erg s-1. While significantly more luminous, the X-ray to UV/optical ratios of the ENTs are broadly similar to TDEs and ANTs as well as within the typical range of AGNs (see the supplementary text), again supporting an accretion-based origin for these events.
@@
The rate of luminous, long-lived, accretion-powered events like these ENTs can also be used to understand their potential physical origins. The high redshift and low observed number of ENTs is immediately a sign of a low intrinsic rate. From the 3 detected ENTs, their peak absolute magnitudes, and the survey parameters for Gaia and ZTF, we estimate a lower limit on the rate of Gpc-3 yr-1. This implies that ENTs are roughly ten thousand times less common than SLSNe and TDEs at (?, ?).
@@
With the observed properties of the ENT flares and their estimated rates, we consider potential physical models and plausible origins. First, we examine strong gravitational lensing, which can significantly magnify transient events (e.g., (?)). The strong constraints on the lack of a foreground lens galaxy, even when considering the effects of magnification bias (?), and the high required magnifications ( for normal SNe) make this a remote possibility. Typical radioactively-powered SNe are ruled out based on the unphysically high ( M⊙) 56Ni masses that they would require. There are several luminous classes of SNe (?), including those powered by magnetar spin-down and interactions with circumstellar material (CSM). A magnetar-powered event is ruled out as the most energetic ENT would require a neutron star spinning at breakup to be 5.5 M⊙, even assuming 100% efficiency. This mass is well above the Tolman–Oppenheimer–Volkoff upper limit on the mass of a neutron star (e.g., (?)). CSM interactions are also ruled out as they predominantly produce narrow emission lines, which are not seen for all of the ENTs, and the CSM masses required are of order M⊙. Thus, no stellar transient can be responsible for the ENTs.
@@
As there is evidence for AGN activity, albeit typically weak, in most of the ENT host galaxies, an AGN origin for the flares must be considered. From studies on quasar variability, a flare resulting from an extreme stochastic variability event is ruled out (?). Another class of transients requiring interaction between an AGN disk wind and the broad line region clouds (?) is unlikely given the high required masses and the sub-Eddington pre-flare accretion rates. Finally, we find it unlikely that instabilities within an AGN disk cause the ENTs due to the similar accreted masses and timescales of the events despite a large range in pre-flare Eddington ratios of the AGNs within the ENT host galaxies.
@@
The most plausible physical scenario for these ENTs is the tidal disruption of an intermediate-mass star and the subsequent return of material onto the SMBH. The high masses of the SMBHs naturally provide long-duration flares consistent with the ENT timescales, as the flare timescale scales as M. As roughly half of the disrupted stellar mass in a TDE leaves the system, the total radiated energies provide a lower limit on the stellar masses of M⊙. The timescales and luminosities for the disruption of M⊙ stars match the ENT observables well. Scaling from the known TDE rate and assuming that the TDE rate is proportional to the number of stars and their stellar lifetimes, we estimate the rate of M⊙ TDEs at to be Gpc-3 yr-1, in good agreement with the estimated ENT rate. As the ENT rate is formally a lower limit, we note that top-heavy stellar initial mass functions (?) and lower disrupted stellar masses can both increase the expected rates. Given the natural explanation of smooth flares, the compatible timescale and luminosities, and several multi-wavelength similarities we propose that the ENTs are the product of the tidal disruption of intermediate-mass stars.

### v2

# The Most Energetic Transients: Tidal Disruptions of High-Mass Stars

@@
The Most Energetic Transients:
Tidal Disruptions of High-Mass Stars
We present the class of extreme nuclear transients (ENTs), including the most energetic single transient yet discovered, Gaia18cdj. Each ENT is coincident with its host-galaxy nucleus and exhibits a smooth (% excess variability), luminous ( to erg s-1), and long-lived ( days) flare. ENTs are extremely rare ( Gpc-3 yr-1) compared to any other known class of transients. They are at least twice as energetic ( to erg) as any other known transient, ruling out supernova origins. Instead, the high peak luminosities, long flare timescales, and immense radiated energies of the ENTs are most consistent with the tidal disruption of high-mass ( M⊙) stars by massive ( M⊙) supermassive black holes (SMBHs). ENTs will be visible to high redshifts ( to ) in upcoming surveys, providing an avenue to study the high-mass end of the SMBH mass distribution, complementing recent studies of actively accreting SMBHs at high redshifts with the James Webb Space Telescope.
Introduction
@@
Accretion onto supermassive black holes (SMBHs) powers many of the most luminous events in the universe. At a redshift of , roughly 10% of SMBHs are actively accreting mass (e.g., (?)) and are observed as active galactic nuclei (AGNs). AGN light curves commonly show stochastic variability at a broad range of timescales from minutes to years (e.g., (?)), with some AGNs showing long-term photometric trends often accompanied by dramatic changes in their spectra (e.g., (?, ?)). AGNs can also, albeit rarely, exhibit large, coherent flares (?), although the physical mechanisms for powering them are unclear.
@@
With the recent growth of optical transient surveys, several classes of flares coincident with the nuclei of their host galaxies have been detected. These include tidal disruption events (TDEs; (?, ?)), rapid turn-on AGNs (?, ?), and ambiguous nuclear transients (ANTs; (?, ?, ?)). Accretion-powered transients share several key observational properties, including bright ultraviolet (UV) emission (?, ?, ?, ?, ?), strong emission lines (?, ?, ?), and often X-ray emission (?, ?, ?, ?). The smooth flares of nuclear transients on several-month timescales (?, ?) are distinct from the stochastic variability typical of AGNs.
@@
A TDE results from the disruption of a star as it passes too close to a SMBH (e.g., (?, ?)). Most observed TDEs appear consistent with the disruption of a main-sequence star with a mass of M⊙ (?), although there is appreciable scatter in these estimates (?). Nevertheless, characteristics like enhanced N/C ratios (?, ?) suggest a population of TDEs resulting from more massive stars. The host galaxies of TDEs typically do not host a strong AGN, although this is likely driven in part by selection effects (?, ?). Recently, an increasing number of TDE candidates have been discovered for which their host galaxies exhibit signs of weak AGN activity (e.g., (?, ?)).
Results
@@
Owing to the establishment of long-baseline all-sky surveys (?, ?, ?), we are now sensitive to rare and unexpected classes of transients. One such survey, Gaia Alerts, uses the Gaia spacecraft to monitor the transient sky at approximately monthly cadence with a per-transit 5 depth of 21 mag. From the Gaia Alerts (?, ?) transient stream we selected a sample of flares with three primary characteristics: (1) large amplitudes of 1 mag, (2) smooth light curves with 10% excess variability about the flare evolution, and (3) a long timescale of year. Gaia is ideal for such a search as it has observed the full sky since late 2014 and, as a space-based mission, it typically has shorter seasonal breaks than ground-based surveys. Our search yielded two transients, Gaia16aaw (AT2016dbs) and Gaia18cdj (AT2018fbb). We combine these events with the recently published object ZTF20abrbeie (AT2021lwx; (?, ?)) as a sample of events we will refer to as extreme nuclear transients (ENTs).
@@
The observed properties of the ENTs are reminiscent of extreme versions of ANTs, which are transients occurring in an AGN host galaxy. The light curves of the ENTs, shown in Figure 1, each exhibit a long ( 100 day) rise to a high peak luminosity. The ENTs decline slowly after peak, taking more than 150 days to fade to half of their peak luminosity. The ENTs detected prior to the flare show tentative signs of pre-flare variability, suggesting weak AGN activity within their host galaxies. After the UV/optical emission peaks, the ENTs show an IR excess, indicative of transient heating of circumnuclear dust and re-emission at longer wavelengths. Much like the ENT hosts, the host galaxies of AGNs typically have large amounts of nuclear dust.
@@
The ENTs Gaia16aaw and Gaia18cdj are located within kpc and kpc of their host-galaxy centers, respectively, confirming that they are nuclear transients. Their nuclear origin, long timescales, and high peak luminosities are immediately suggestive of a transient resulting from accretion onto a SMBH. We cannot measure the host offset of AT2021lwx as it has no detected host galaxy prior to the flare, at levels of absolute mag in the rest-frame blue bands. Nevertheless, the similar timescales and peak luminosities of AT2021lwx indicate that it is also powered by accretion, consistent with previous studies.
@@
The ENTs are located at a relatively high redshift of , as measured from optical and near-IR follow-up spectra. Using the stellar population synthesis and AGN model of the Code Investigating GALaxy Emission (CIGALE; (?)), we find that the host galaxies of Gaia16aaw and Gaia18cdj each have a stellar mass of M⊙ and star formation rates (SFRs) of M⊙ yr-1. While undetected, the luminosity limits for the host galaxy of AT2021lwx combined with a conservative mass-to-light ratio of 3 yields a mass upper-limit of M⊙. The upper limit on the [O II] emission means that the SFR for AT2021lwx is M⊙ yr-1.
@@
From typical galaxy-SMBH scaling relations, with a scatter of 0.4 dex (?), the stellar masses imply SMBH masses of M⊙ for Gaia16aaw and Gaia18cdj and a 3 upper limit on the mass of M⊙ for AT2021lwx, which are more massive SMBHs than those in the majority of known nuclear transient hosts. The detected ENT hosts are more massive and display higher SFRs than the host galaxies of local nuclear transients such as TDEs and ANTs. Furthermore, the detected ENT host-galaxy masses are within the top few percent of stellar masses at , when the universe was half its current age. Thus, the inferred SMBH masses are similarly extreme at this redshift. In contrast, the prodigious SFRs of the hosts of Gaia16aaw and Gaia18cdj are only moderately high, in the 70th percentile of specific star formation rate at , when the star formation density in the universe was a factor of 6 higher than today (?).
@@
The rest-frame spectra of the ENTs, shown in Figure 2, exhibit blue spectra with broad lines from the Balmer series of hydrogen and singly-ionized magnesium (Mg II). These are similar to the comparison spectra of the luminous nuclear transients AT2019brs (?), ASASSN-17jz (?), PS1-10adi (?), ASASSN-18jd (?), PS16dtm (?), and AT2019dsg (?) and broadly similar to the energetic SLSN-II SN2018lzi (?) all shown in Fig. 2. ASASSN-15lh (?, ?) also shows a blue continuum but does not have strong emission lines. The spectrum of the energetic SLSN-I SN2020qlb (?) is distinct from the ENT spectra. The persistent blue continua and broad lines are inconsistent with known classes of supernovae, but fully consistent with SMBH accretion. The Mg II and H emission of the ENTs is broad ( km s-1) and luminous ( erg s-1), very similar to AGNs. While Mg II emission has not been seen for TDEs, the H emission is consistent with TDEs if the lines are equally as overluminous as the broadband emission.
@@
The ENT spectral energy distributions (SEDs) are well-fit by a blackbody model, which is consistent with the super-Eddington accretion expected from TDEs and some supernovae (SNe). In contrast, AGNs with broad lines typically exhibit power-law-like SEDs as a result of viewing the accretion disk directly. The resulting bolometric light curves are shown in Figure 3, along with the comparison objects from Fig. 2. The ENTs are extremely luminous, with peak luminosities of erg s-1. This is times more luminous than typical core-collapse SNe, times the average Type Ia SN peak luminosity, times more luminous than the median Type I superluminous supernovae (SLSN-I; (?)), and times more luminous than the average SLSNe-II (?). The most energetic SLSN-I (SN2020qlb) and SLSN-II (SN2018lzi), shown in Figure 3, are % as luminous as the ENTs and an order of magnitude less energetic. Only ASASSN-15lh, suggested to either be SLSN-I (?)) or TDE (?), rivals these high peak luminosities.
@@
The blackbody properties of the ENTs are also consistent with some form of accretion onto a SMBH rather than an exotic class of SN. The temperatures of the ENTs are hot at K and show little or very slow evolution during the flare. This is inconsistent with SLSNe, which quickly cool as the ejecta expands. However, this behavior is similar to TDEs and ANTs, with a remarkable agreement between the ANT and ENT blackbody temperatures. The large effective radii of the ENTs are consistent with SLSNe and some ANTs, although the decreasing blackbody radii in time are more typical of TDEs and ANTs.
@@
The light curve decay timescale of the ENTs is also far longer than most transients, which is consistent with TDEs occurring on massive SMBHs. The rest-frame durations for the flares to fade by half are days for Gaia16aaw, days for Gaia18cdj, and days for AT2021lwx. Figure 4 shows the position of these ENTs in the parameter space of peak absolute magnitude and characteristic timescale. The ENTs stand out in this parameter space for being very luminous and long-lived. ASASSN-15lh, which has a similar peak luminosity to the ENTs, decays more quickly, with a timescale of days. The position of the ENTs in the upper right corner of this space indicates a high total radiated energy.
@@
The radiated energies of the ENTs are erg for Gaia16aaw, erg for Gaia18cdj, and erg for AT2021lwx. These extreme radiated energies correspond to high accreted masses of M⊙ for a typical 10% accretion efficiency, far greater than typical TDEs and ANTs. The ENT flares are at least twice as energetic as the next most energetic known flares, PS1-10adi (?) and ASASSN-15lh (?, ?), and up to an order of magnitude higher in the cases of Gaia18cdj and AT2021lwx. While the estimates for the accreted mass are extreme compared to local transients, they are fully consistent with the tidal disruption of a high-mass (3 M⊙) star.
@@
The dust properties of ENTs are very similar to other SMBH accretion-powered transients. To study the environment directly surrounding the SMBHs, we used NEOWISE data to probe the emission from hot dust as it reprocesses the intense UV/optical emission from the flare. By fitting the WISE IR SEDs as blackbodies, we find peak luminosities of erg s-1, temperatures of K and radii of pc, all consistent with hot dust in nuclear environments (?, ?). From the ratio of the peak IR luminosity to the peak UV/optical luminosity we estimate dust covering fractions of , which we confirm with models of the optical and IR light curves. These covering fractions are consistent with AGN dust covering fractions at similar SMBH masses as well as dust-obscured TDE candidates (?). This indicates the presence of dense gas and dust near the SMBH, which likely supports the existence of AGN activity, whether weak or in the past, in each of the ENT hosts.
@@
Accretion-powered transients often show X-ray emission, as do two of the three ENTs in our sample. Gaia16aaw and AT2021lwx both show X-rays at levels of erg s-1 in the rest-frame 0.3–10 keV band throughout the flare, similar to luminous AGNs (?). Gaia18cdj is undetected in the X-rays at erg s-1. While significantly more luminous, the X-ray to UV/optical ratios of the ENTs are broadly similar to TDEs and ANTs as well as within the typical range of AGNs, again supporting an accretion-based origin for these events.
As the ENTs are likely powered by accretion, it is important to consider the presence of previous AGN activity. Through a combination of the CIGALE fits, WISE IR colors, narrow emission lines, and X-ray emission we find evidence of a strong AGN in Gaia16aaw. Gaia18cdj likely hosts a weak AGN based on an [O III] line luminosity similar to Seyferts, but the CIGALE fits and MIR colors rule out a strong quasar that dominates the observed emission. The rest-frame UV luminosity and [O III] line luminosity limits derived from the pre-flare properties of AT2021lwx rule out a strong quasar but remain consistent with a weaker AGN. The fact that some ENT host galaxies do not host a strong AGN suggests that prior strong AGN activity is not a requirement to power an ENT.
@@
The rate of luminous, long-lived, accretion-powered events like these ENTs can also be used to understand their potential physical origins. The high redshift and low observed number of ENTs is suggestive of a low intrinsic rate. From the 3 detected ENTs, their peak absolute magnitudes, and the survey parameters for Gaia and ZTF, we estimate a lower limit on the rate of Gpc-3 yr-1. Further examination of selection effects and biases in searching for these extreme transients is needed to refine estimates of the intrinsic rate of ENTs. Nevertheless, our estimated rate implies that ENTs are roughly ten thousand times less common than SLSNe and TDEs at (?, ?).
Discussion
@@
With the observed properties of the ENT flares and their estimated rates, we consider potential physical models and plausible origins. First, we examine strong gravitational lensing, which can magnify transient events (e.g., (?)). Constraints on the lack of a foreground lens galaxy from photometry and spectra, even when considering the effects of magnification bias (?), and the high required magnifications ( for normal SNe) make this a remote possibility. Typical radioactively-powered SNe are ruled out based on the unphysically high ( M⊙) 56Ni masses that they would require. There are several luminous classes of SNe (?), including those powered by magnetar spin-down and interactions with circumstellar material (CSM). A magnetar-powered event is ruled out as the most energetic ENT would require a neutron star spinning at breakup to be 5.5 M⊙, even assuming 100% efficiency. This mass is well above the Tolman–Oppenheimer–Volkoff upper limit on the mass of a neutron star (e.g., (?)). CSM interactions are also ruled out as they predominantly produce narrow emission lines, which are not seen for all of the ENTs, and the required CSM masses are of order M⊙. Thus, no stellar transient can be responsible for the ENTs.
@@
As there is evidence for AGN activity, albeit typically weak, in most of the ENT host galaxies, an AGN origin for the flares must be considered. From studies on quasar variability, a flare resulting from an extreme stochastic variability event is ruled out (?) given estimated rates of smooth AGN flares an order of magnitude below the ENT rate. Another class of transients requiring interaction between an AGN disk wind and the broad line region clouds (?) is unlikely given the high required masses and the sub-Eddington pre-flare accretion rates. Finally, we find it unlikely that instabilities within an AGN disk cause the ENTs due to the similar accreted masses and timescales of the events despite a large range in pre-flare Eddington ratios of the AGNs within the ENT host galaxies, although they cannot be ruled out entirely.
@@
The most plausible physical scenario for these ENTs is the tidal disruption of a high-mass star and the subsequent return of material onto the SMBH. The high masses of the SMBHs naturally provide long-duration flares consistent with the ENT timescales, as the flare timescale scales as . Thus, for a TDE occurring on the M⊙ SMBHs of the ENT hosts, we would expect a timescale at least 5 times longer than typical TDEs occurring on M⊙ SMBHs. As roughly half of the disrupted stellar mass in a TDE leaves the system, the total radiated energies provide a lower limit on the stellar masses of M⊙. The timescales and luminosities for the disruption of M⊙ stars match the ENT observables well (?, ?). Scaling from the known local TDE rate and assuming that the TDE rate is proportional to the number of stars, using an IMF with high-mass slope of , and their stellar lifetimes we estimate the rate of M⊙ TDEs at to be Gpc-3 yr-1, consistent with the estimated ENT rate. As the ENT rate is formally a lower limit, we note that top-heavy stellar initial mass functions (?) can increase the expected rates. We additionally find several plausible explanations for the difference in observed rates and the theoretical estimate above. Each ENT exhibits a large dust echo, which implies that a large fraction might be obscured and therefore missing from optical surveys. Additionally, as Gaia did not trigger on AT2021lwx, these surveys are incomplete. Both of these effects will result in a higher intrinsic ENT rate than our estimate. Finally, the theoretical rate estimate is weighted towards the lowest mass stars. The theoretically expected rates would decrease if the ENTs are powered by stars more massive than our assumed M⊙ lower bound.
Given the natural explanation of smooth flares (?, ?), the compatible timescale and luminosities, and several multi-wavelength similarities we propose that the ENTs are the product of the tidal disruption of high-mass stars. The presence of AGNs in the ENT host galaxies is likely related to the order of magnitude increase in the AGN fraction at as compared to local galaxies (?). Additionally, recent work suggests that AGN disks may increase TDE rates substantially (?, ?). Both effects make it more likely that TDEs at high redshifts will occur in AGN host galaxies.